Problem ellipsomter SNF 2007-09-25 18:56:49: Laser beam problem

khanh at snf.stanford.edu khanh at snf.stanford.edu
Tue Sep 25 18:56:50 PDT 2007


I noticed that the beam on the chuck on and off and seem like the beam is blocked sometimes.
Measured LTO oxide on si (expected 200) and gave me 4 different values at the same spot on the wafer 150, 201, 750 and  1000.




More information about the ellipsomter-pcs mailing list